sporting store sold twice as many t-shirts as shorts. T-shirts are 9.00 each, and shorts are 14.00 each. the total amunt of money taken in for both items was 256.00 find the number for each that was sold

Answer by checkley77(12844)

t=2s
9t+14s=256
9*2s+14s=256
18s+14s=256
32s=256
s=256/32
s=$8.00 for each pair of shorts
t=2*8=$16.00 for each t-shirt.
Proof
9*16+14*8=256
144+112=256
256=256